A boy who is riding a bicycle moves with an initial velocity of 5 m /s. Ten

seconds later, he is moving at 15 m/s. what is his acceleration?


1
Expert's answer
2022-02-18T15:53:39-0500

Given:

"v_0=5\: \rm m/s"

"v=15\: \rm m/s"

"\Delta t=10\:\rm s"


The aceeleration of a bicycle

"a=\frac{\Delta v}{\Delta t}=\frac{15-5}{10}=1\:\rm m/s^2"